OpenStudy (anonymous):

A savvy shopper combines a 25% off coupon with the store’s existing 25% off sale. Ignoring sales tax, how much of the base price should the shopper expect to pay? a. 75% b. 56.25% c. 50% d. 43.75%

OpenStudy (anonymous):

I think it is B. 56.25% Ex: 100 * 75% (what you'll pay) = 75*75% = 56.25

OpenStudy (anonymous):

0.75 (0.75) = 56.25%

OpenStudy (anonymous):

thank u